Perdon Corporation manufactures safes—large mobile safes, and large walk-in stationary bank safes. As part of its annual budgeting process, Perdon is analyzing the profitability of its two products. Part of this analysis involves estimating the amount of overhead to be allocated to each product line. The information shown below relates to overhead.

Mobile Safes

Walk-in Safes

Units planned for production 200 50 Material moves per product line 300 200 Purchase orders per product line 450 350 Direct labor hours per product line 800 1,700

The total estimated manufacturing overhead was $276,000. Under traditional costing (which assigns overhead on the basis of direct labor hours), what amount of manufacturing overhead costs are assigned to: (Round answers to 2 decimal places, e.g. 12.25.)

The total estimated manufacturing overhead of $276,000 was comprised of $172,000 for materials handling costs and $104,000 for purchasing activity costs. Under activity-based costing (ABC): (Round answers to 2 decimal places, e.g. 12.25.)

What amount of materials handling costs are assigned to:
one mobil safe = 516

one walk in safe = 1376

The total estimated manufacturing overhead of $276,000 was comprised of $172,000 for materials handling costs and $104,000 for purchasing activity costs. Under activity-based costing (ABC): (Round answers to 2 decimal places, e.g. 12.25.)

What amount of purchasing activity costs are assigned to:

one mobil safe per unit =

one walk in safe per unit =
Answer #1

Purchasing activity rate = 104000/800 = 130 per purchase order

One mobile safe per unit = (130*450)/200 = 292.50

One walk in safe per unit = (130*350)/50 = 910